# Model Merging in the Era of Large Language Models: Methods, Applications, and Future Directions

### 10 Mar 2026 v1

@@ 2.3 Weight Space Symmetries and Permutation Invariance
These fundamental symmetries, particularly permutation invariance, substantially impact the feasibility and methodology of model merging. Permutation invariance arises from that reordering hidden units within a layer, along with corresponding adjustments to incoming and outgoing weight matrices, produces a functionally equivalent network. This mathematical property implies that the number of distinct parameterizations encoding identical input-output mappings grows combinatorially with network depth and width, creating substantial challenges for direct weight space operations.
@@
Consider a feedforward network with weight matrices $W^{(l)}$ connecting layer l1l-1 to layer $l$. For any permutation matrix $P^{(l)}$ applied to the hidden units of layer $l$, the transformation $(W^{(l)},W^{(l+1)})\mapsto(P^{(l)}W^{(l)},W^{(l+1)}(P^{(l)})^{T})$ preserves the network’s function exactly. For a network with $L$ hidden layers each containing $n$ units, this yields $(n!)^{L}$ equivalent parameterizations, a symmetry group whose size renders the loss landscape highly non-convex despite potential functional simplicity. The comprehensive treatment in (Ainsworth et al., 2023) emphasizes that this symmetry significantly complicates the interpretation of distances and interpolations in weight space, as two networks computing identical functions may appear arbitrarily distant under Euclidean metrics.
@@
The implications for model merging are immediate and severe. When averaging weights from two independently trained networks, the hidden unit orderings are essentially arbitrary and unrelated, causing the averaged network to combine misaligned features in semantically meaningless ways. Formally, if $\theta_{A}$ and $\theta_{B}$ represent two solutions related by permutation such that $\theta_{B}=\pi(\theta_{A})$ for some permutation $\pi$, then $\frac{1}{2}(\theta_{A}+\theta_{B})$ generally lies far from both solutions in functional space despite occupying their geometric midpoint. This observation explains why naive weight averaging of independently trained networks typically produces catastrophic performance degradation, even when both source networks achieve comparable accuracy.
@@
The theoretical framework connecting permutation alignment (Ainsworth et al., 2023) to optimal transport provides principled methods for resolving this challenge. Given two networks with weight matrices $W_{A}$ and $W_{B}$, finding the optimal permutation reduces to solving an assignment problem that minimizes some distance metric between aligned representations. Ferbach et al. (2024) discuss how this formulation enables tractable algorithms based on the Hungarian method or approximate transport solvers, establishing correspondence between hidden units based on their functional roles rather than arbitrary indexing.
@@
The work of Nobari et al. (2025) demonstrated that even after permutation alignment, residual distribution mismatches in intermediate activations can create apparent barriers to interpolation. Their analysis revealed that permutation-aligned networks may still exhibit incompatible feature statistics, means and variances of hidden activations, that cause interpolated networks to produce out-of-distribution internal representations. This insight motivates post-alignment normalization procedures that reconcile activation distributions, further reducing barriers to successful merging.
@@
Shared initialization provides a natural mechanism for symmetry breaking that circumvents permutation ambiguity entirely. When multiple models are fine-tuned from identical pretrained weights, the optimization trajectories implicitly preserve hidden unit correspondence established during pretraining. This explains the empirical observation that fine-tuned models exhibit substantially better mergeability than independently trained counterparts (Wortsman et al., 2022a). The pretrained initialization effectively anchors each hidden unit to a specific functional role, ensuring that corresponding parameters across fine-tuned variants encode semantically related modifications.
@@
Recent systematic investigations have quantified the relationship between alignment quality and merging success, revealing that architectural choices and training procedures influence the degree to which permutation alignment is necessary. Transformer architectures with their structured attention mechanisms exhibit different symmetry properties than fully-connected networks. These insights regarding weight space symmetries, combined with the earlier discussions of loss landscape geometry and mode connectivity, collectively inform a unified understanding of the conditions necessary for successful model merging.
@@ 2.4 Prerequisites and Conditions for Successful Model Merging
This unified understanding can be formalized as a framework of mergeability conditions, the set of prerequisites that must be satisfied for model merging to yield high-quality results. Understanding these conditions is essential for practitioners seeking to predict whether a given set of models can be successfully combined and for researchers developing new merging algorithms with broader applicability.
@@
The most critical prerequisite for successful model merging is shared pretrained initialization, which establishes a common reference point in weight space from which all candidate models diverge during fine-tuning. Formally, consider a pretrained model with parameters $\theta_{\text{pre}}$ and a set of fine-tuned models $\{\theta_{1},\theta_{2},\ldots,\theta_{K}\}$ each obtained by optimizing task-specific objectives starting from $\theta_{\text{pre}}$. The shared initialization ensures that fine-tuned parameters remain within the same loss basin, enabling linear interpolation $\theta_{\alpha}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}$ to traverse low-loss regions rather than crossing high-loss barriers (Garipov et al., 2018; Frankle et al., 2020). This basin-sharing property significantly distinguishes the fine-tuning paradigm from independent training, where models converge to disconnected minima separated by substantial energy barriers.
@@
Architectural compatibility is a second essential condition, requiring both structural identity and representational alignment. Models must share identical architectures—layer configurations, activation functions, and normalization schemes—for element-wise parameter operations to be well-defined. Beyond this structural requirement, Li et al. (2023a) demonstrate that certain architectural choices exhibit inherently favorable merging properties. Transformer architectures (Vaswani et al., 2017), with their residual connections and layer normalization, maintain more consistent feature representations across fine-tuning variants, facilitating parameter combination without the severe interference observed in earlier network designs. The residual stream structure effectively constrains the magnitude of per-layer modifications, keeping fine-tuned models closer to the pretrained initialization.
@@
Training procedure similarity also strongly influences merging success. Models fine-tuned with similar hyperparameters (learning rate, batch size, optimizer) tend to follow comparable trajectories through weight space, yielding stronger linear mode connectivity (Singh and Jaggi, 2020). Conversely, substantial procedural divergence can cause models to settle in distinct sub-basins within the broader pretrained basin, introducing interpolation barriers that degrade merge quality. Izmailov et al. (2018) provide empirical analysis showing how training duration and learning rate magnitude influence the distance between fine-tuned solutions and consequently the loss barrier encountered during interpolation.
@@
The theoretical framework for predicting capability preservation during merging remains incomplete but can be partially characterized through the lens of task vector interference. When fine-tuning induces parameter modifications in complementary subspaces, different parameters or opposing directions for different tasks, merging preserves individual capabilities through linear superposition. However, when task-specific adaptations overlap substantially and conflict in direction, the merged model exhibits degraded performance on all constituent tasks. Nobari et al. (2025) demonstrate that even when parameter-level conflicts are resolved, activation distribution mismatches can cause capability degradation, highlighting the multi-faceted nature of mergeability assessment. This observation motivates the development of mergeability prediction methods that analyze both parameter-space and representation-space compatibility, enabling principled decisions about which models can be successfully combined. Yet while these conditions provide practical guidance, the theoretical foundations underlying them, particularly for large-scale language models, remain incompletely understood.
@@ 2.5 Open Theoretical Questions and Research Frontiers
Despite this progress, several substantial gaps persist between observed phenomena and rigorous theoretical explanations, presenting fertile ground for future research. The most pressing open question concerns the fundamental mechanisms by which pretrained large language models exhibit such strong linear mode connectivity (Singh and Jaggi, 2020) following independent fine-tuning on disparate tasks. While it is well-established that shared initialization creates a common loss basin, the theoretical understanding of why this basin remains sufficiently broad and convex to accommodate diverse task-specific adaptations, particularly as model scale increases to billions of parameters, remains incomplete. (Singh et al., 2024) identifies this explanatory gap as a central challenge, noting that existing mode connectivity theory was largely developed in smaller neural networks and may require substantial extension to account for the unique properties of large-scale language models.
@@
The relationship between model scale and mergeability properties presents another theoretically unresolved frontier. Empirical observations suggest that larger models exhibit more favorable merging characteristics, with reduced loss barriers and improved capability retention compared to smaller counterparts. However, a principled mathematical framework explaining this phenomenon remains elusive. One hypothesis posits that increased overparameterization expands the effective dimensionality of the loss basin, providing more ”room” for task-specific adaptations to coexist without interference. An alternative perspective suggests that the emergent representations in larger models may be more compositional and modular, naturally supporting the superposition of task-specific modifications. Recent work provides empirical investigation into scale effects but acknowledges that theoretical explanations for observed scaling behaviors remain speculative.
@@
Emerging research directions seek to develop predictive frameworks for merge compatibility that could obviate the need for exhaustive experimentation. The challenge can be formalized as learning a function $\mathcal{C}:(\theta_{1},\theta_{2},\ldots,\theta_{k})\rightarrow[0,1]$ that estimates the expected performance of merging models $\theta_{1},\ldots,\theta_{k}$ without actually performing the merge. Promising approaches analyze task vector statistics, including angular relationships and magnitude distributions, to predict interference severity. surveys existing techniques for characterizing solution relationships through loss landscape analysis, providing foundational tools that could support compatibility prediction. However, translating these geometric insights into reliable predictive models requires addressing the computational challenge of efficiently approximating high-dimensional loss surface properties.
@@
Theoretical investigations into the fundamental limits of merging constitute another critical research frontier. While sparsification-based methods such as TIES-Merging and DARE have demonstrated that parameter interference can be mitigated, the question of when interference becomes irreconcilable, representing a fundamental rather than technical limitation, remains open. Nobari et al. (2025) demonstrate that activation-level statistics provide complementary information to parameter-level analysis, suggesting that comprehensive mergeability theory must account for both weight space and representation space phenomena. Furthermore, the conditions under which nonlinear interpolation paths offer strictly superior outcomes compared to linear combinations, and how such paths might be efficiently discovered, represent underexplored theoretical territories with significant practical implications.
@@
Finally, extending merging theory to accommodate architectural heterogeneity presents significant theoretical challenges. Current frameworks assume exact architectural correspondence, but practical scenarios increasingly involve models with variations in layer counts, hidden dimensions, or attention mechanisms. Developing principled theoretical foundations for cross-architecture merging, potentially leveraging representation alignment or functional correspondence rather than direct parameter mapping, could substantially expand the applicability of model merging techniques while requiring novel mathematical frameworks that transcend current weight-space formulations.
Table 2: Comparison of theoretical foundations and concepts in model merging.
@@
Loss Landscape Geometry Analyzes geometric properties of high-dimensional parameter space Explains why weight interpolation yields functional networks Complex landscapes make complete characterization difficult 2024
@@
(Garipov et al., 2018)
Identifies low-loss paths between trained solutions in weight space Central to explaining success of weight interpolation methods Requires shared initialization for strong connectivity 2024
@@
Exploits symmetry from reordering hidden units with weight adjustments Multiple equivalent parameterizations for same function Alignment required before merging independently trained models 2024
@@
Analyzes fundamental symmetry properties affecting merging feasibility Enables functionally equivalent network representations Symmetry breaking needed for meaningful comparisons 2024
@@
(Singh and Jaggi, 2020)
Explains role of common pretrained weights in enabling merging Strong predictor of linear mode connectivity Theoretical mechanisms not fully understood for LLMs 2024
Convexity Analysis Examines local and global convexity properties of loss surfaces Convex regions enable reliable weight averaging Neural networks generally non-convex globally 2024
The theoretical foundations examined in this section, encompassing loss landscape geometry, linear mode connectivity (Singh and Jaggi, 2020), and weight space symmetries, collectively establish the mathematical conditions under which model merging can succeed (summarized in Table 2), while also highlighting the fundamental barriers that arise when models reside in disconnected basins or exhibit misaligned permutation symmetries. These insights provide essential guidance for understanding when simple averaging will preserve model capabilities and when more sophisticated interventions become necessary. With this theoretical groundwork established, the following section turns to practical weight-space averaging and geometric interpolation methods, which translate these foundational principles into the algorithmic methods that instantiate these principles.

### 30 Mar 2026 v2

@@ 2.3 Weight Space Symmetries and Permutation Invariance
These fundamental symmetries, particularly permutation invariance, substantially impact the feasibility and methodology of model merging. Permutation invariance arises from the fact that reordering hidden units within a layer, along with corresponding adjustments to incoming and outgoing weight matrices, produces a functionally equivalent network. This mathematical property implies that the number of distinct parameterizations encoding identical input-output mappings grows combinatorially with network depth and width, creating major challenges for direct weight space operations.
@@
Consider a feedforward network with weight matrices $W^{(l)}$ connecting layer l1l-1 to layer $l$. For any permutation matrix $P^{(l)}$ applied to the hidden units of layer $l$, the transformation $(W^{(l)},W^{(l+1)})\mapsto(P^{(l)}W^{(l)},W^{(l+1)}(P^{(l)})^{T})$ preserves the network’s function exactly. For a network with $L$ hidden layers each containing $n$ units, this yields $(n!)^{L}$ equivalent parameterizations, a symmetry group whose size renders the loss surface highly non-convex despite potential functional simplicity. The detailed treatment in (Ainsworth et al., 2023) emphasizes that this symmetry greatly complicates the interpretation of distances and interpolations in weight space, as two networks computing identical functions may appear arbitrarily distant under Euclidean metrics.
@@
For model merging, the implications are immediate and severe. When averaging weights from two independently trained networks, the hidden unit orderings are essentially arbitrary and unrelated, causing the averaged network to combine misaligned features in semantically meaningless ways. Formally, if $\theta_{A}$ and $\theta_{B}$ represent two solutions related by permutation such that $\theta_{B}=\pi(\theta_{A})$ for some permutation $\pi$, then $\frac{1}{2}(\theta_{A}+\theta_{B})$ generally lies far from both solutions in functional space despite occupying their geometric midpoint. This observation explains why naive weight averaging of independently trained networks typically produces catastrophic performance degradation, even when both source networks achieve comparable accuracy.
@@
A theoretical framework connecting permutation alignment (Ainsworth et al., 2023) to optimal transport provides rigorous methods for resolving this challenge. Given two networks with weight matrices $W_{A}$ and $W_{B}$, finding the optimal permutation reduces to solving an assignment problem that minimizes some distance metric between aligned representations. Practical implementations use the Hungarian method or approximate transport solvers to establish correspondence between hidden units based on their functional roles rather than arbitrary indexing. In a complementary line of work, Ferbach et al. (2024) prove that sufficiently wide two-layer networks trained with SGD are linearly mode connected with high probability, using convergence rates of empirical measures in Wasserstein distance; this provides rigorous theoretical support for the empirical success of permutation-based alignment methods.
@@
The REPAIR framework (Jordan et al., 2023) showed that even after permutation alignment, residual distribution mismatches in intermediate activations can create apparent barriers to interpolation. Permutation-aligned networks may still exhibit incompatible feature statistics that cause interpolated networks to produce out-of-distribution internal representations. This insight motivates post-alignment normalization procedures that reconcile activation distributions, further reducing barriers to successful merging.
@@
Shared initialization provides a natural mechanism for symmetry breaking that circumvents permutation ambiguity entirely. When multiple models are fine-tuned from identical pretrained weights, the optimization trajectories implicitly preserve hidden unit correspondence established during pretraining. This explains the empirical observation that fine-tuned models exhibit considerably better mergeability than independently trained counterparts (Wortsman et al., 2022a). The pretrained initialization effectively anchors each hidden unit to a specific functional role, ensuring that corresponding parameters across fine-tuned variants encode semantically related modifications.
@@
Recent systematic investigations have quantified the relationship between alignment quality and merging success, revealing that architectural choices and training procedures influence the degree to which permutation alignment is necessary. Transformer architectures with their structured attention mechanisms exhibit different symmetry properties than fully-connected networks. Together with the loss surface geometry and mode connectivity results of the preceding subsections, these symmetry considerations form a complete theoretical picture, namely that merging succeeds when models share a loss basin (geometry), can be connected by low-loss paths (connectivity), and have their hidden units aligned (symmetry). The practical question becomes how to verify and enforce these conditions, which we formalize next as mergeability prerequisites.
@@ 2.4 Prerequisites and Conditions for Successful Model Merging
The theoretical pillars above can be consolidated into a practical framework of mergeability conditions, i.e., the set of prerequisites that must be satisfied for model merging to yield high-quality results. Understanding these conditions is essential for practitioners seeking to predict whether a given set of models can be successfully combined and for researchers developing new merging algorithms with broader applicability.
@@
Perhaps the most critical prerequisite for successful model merging is shared pretrained initialization, which establishes a common reference point in weight space from which all candidate models diverge during fine-tuning. Formally, consider a pretrained model with parameters $\theta_{\text{pre}}$ and a set of fine-tuned models $\{\theta_{1},\theta_{2},\ldots,\theta_{K}\}$ each obtained by optimizing task-specific objectives starting from $\theta_{\text{pre}}$. The shared initialization ensures that fine-tuned parameters remain within the same loss basin, so that linear interpolation $\theta_{\alpha}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}$ with $\sum_{k}\alpha_{k}=1$ traverses low-loss regions rather than crossing high-loss barriers (Garipov et al., 2018; Frankle et al., 2020). This basin-sharing property sharply distinguishes the fine-tuning regime from independent training, where models converge to disconnected minima separated by substantial energy barriers.
@@
Architectural compatibility is a second essential condition, requiring both structural identity and representational alignment. Models must share identical architectures (layer configurations, activation functions, and normalization schemes) for element-wise parameter operations to be well-defined. Transformer architectures (Vaswani et al., 2017), with their residual connections and layer normalization, maintain more consistent feature representations across fine-tuning variants, enabling parameter combination without the severe interference observed in earlier network designs. The residual stream structure effectively constrains the magnitude of per-layer modifications, keeping fine-tuned models closer to the pretrained initialization.
@@
Training procedure similarity also strongly influences merging success. Models fine-tuned with similar hyperparameters (learning rate, batch size, optimizer) tend to follow comparable trajectories through weight space, yielding stronger linear mode connectivity (Frankle et al., 2020). Conversely, marked procedural divergence can cause models to settle in distinct sub-basins within the broader pretrained basin, introducing interpolation barriers that degrade merge quality.
@@
A complete theoretical framework for predicting capability preservation during merging remains elusive but can be partially characterized through the lens of task vector interference. When fine-tuning induces parameter modifications in complementary subspaces, different parameters or opposing directions for different tasks, merging preserves individual capabilities through linear superposition. However, when task-specific adaptations overlap appreciably and conflict in direction, the merged model exhibits degraded performance on all constituent tasks. Activation-informed approaches (Nobari et al., 2025) further show that incorporating activation-level information into the merging process can help preserve capabilities by selectively prioritizing critical weights, highlighting that mergeability assessment should account for both parameter-space and representation-space factors. This observation motivates the development of mergeability prediction methods that analyze both parameter-space and representation-space compatibility, supporting informed decisions about which models can be successfully combined. Yet while these conditions provide actionable guidance, the theoretical foundations underlying them remain incompletely understood, particularly for large-scale language models with billions of parameters.
@@ 2.5 Open Theoretical Questions and Research Frontiers
Several important gaps persist between observed merging phenomena and rigorous theoretical explanations, presenting fertile ground for future research. The most pressing open question concerns the mechanisms by which pretrained large language models exhibit such strong linear mode connectivity (Frankle et al., 2020) following independent fine-tuning on disparate tasks. While shared initialization creates a common loss basin, our understanding of why this basin remains sufficiently broad and convex to accommodate diverse task-specific adaptations remains incomplete, particularly as model scale increases to billions of parameters. Singh et al. (2024) analyze the topological conditions under which linear mode connectivity arises, but extending such analysis to the scale and complexity of modern LLMs remains an open challenge.
@@
How model scale relates to mergeability properties presents another theoretically unresolved frontier. Empirical observations suggest that larger models exhibit more favorable merging characteristics, with reduced loss barriers and improved capability retention compared to smaller counterparts. However, a formal mathematical account of this phenomenon remains elusive. One hypothesis posits that increased overparameterization expands the effective dimensionality of the loss basin, providing more ”room” for task-specific adaptations to coexist without interference. An alternative perspective suggests that the emergent representations in larger models may be more compositional and modular, naturally supporting the superposition of task-specific modifications. Resolving these hypotheses through controlled scaling experiments remains an important open direction.
@@
Emerging research directions seek to develop predictive frameworks for merge compatibility that could obviate the need for exhaustive experimentation. The challenge can be formalized as learning a function $\mathcal{C}:(\theta_{1},\theta_{2},\ldots,\theta_{k})\rightarrow[0,1]$ that estimates the expected performance of merging models $\theta_{1},\ldots,\theta_{k}$ without actually performing the merge. Promising approaches analyze task vector statistics, including angular relationships and magnitude distributions, to predict interference severity. However, translating these geometric insights into reliable predictive models requires overcoming the computational challenge of efficiently approximating high-dimensional loss surface properties.
@@
Theoretical investigations into the fundamental limits of merging constitute another critical research frontier. While sparsification-based methods such as TIES-Merging and DARE have shown that parameter interference can be mitigated, the question of when interference becomes irreconcilable, representing a fundamental rather than technical limitation, remains open. Nobari et al. (2025) show that activation-level statistics provide complementary information to parameter-level analysis, suggesting that a complete mergeability theory must account for both weight space and representation space phenomena. In addition, the conditions under which nonlinear interpolation paths offer strictly superior outcomes compared to linear combinations, and how such paths might be efficiently discovered, represent underexplored theoretical territories with large practical implications.
@@
Finally, extending merging theory to accommodate architectural heterogeneity presents formidable theoretical challenges. Current frameworks assume exact architectural correspondence, but practical scenarios increasingly involve models with variations in layer counts, hidden dimensions, or attention mechanisms. Developing rigorous theoretical foundations for cross-architecture merging, potentially using representation alignment or functional correspondence rather than direct parameter mapping, could substantially expand the applicability of model merging techniques while requiring novel mathematical frameworks that transcend current weight-space formulations.
Table 2: Theoretical foundations underlying model merging success. Each row summarizes a key theoretical concept, its mathematical basis, its implications for merging, and relevant seminal references. Together, these concepts explain when and why weight-space combination of fine-tuned models preserves or enhances performance.
@@
Loss Surface Geometry
(Singh et al., 2024)
Analyzes geometric properties of high-dimensional parameter space Explains why weight interpolation yields functional networks Complex landscapes make complete characterization difficult 2024
@@
(Frankle et al., 2020)
Establishes that models from shared initialization are connected by low-loss linear paths Central to explaining success of weight interpolation methods Requires shared initialization for strong connectivity 2020
@@
Exploits symmetry from reordering hidden units with weight adjustments Multiple equivalent parameterizations for same function Alignment required before merging independently trained models 2023
@@
Analyzes fundamental symmetry properties affecting merging feasibility Enables functionally equivalent network representations Symmetry breaking needed for meaningful comparisons 2023
@@
(Frankle et al., 2020)
Explains role of common pretrained weights in enabling merging Strong predictor of linear mode connectivity Theoretical mechanisms not fully understood for LLMs 2020
Taken together, the theoretical foundations examined in this section, encompassing loss surface geometry, linear mode connectivity (Frankle et al., 2020), and weight space symmetries, collectively establish the mathematical conditions under which model merging can succeed (summarized in Table 2). These insights guide understanding of when simple averaging will preserve model capabilities and when more targeted interventions become necessary. Building on these theoretical principles, the next section examines how they translate into practical merging algorithms.